In Young's double slit experiment, the phase

difference between two interfering beams is

(pi/2). If the wavelength of light is 580 nm,

calculate the path difference between the

interfering beams.


1
Expert's answer
2021-06-14T08:18:20-0400

"\Delta=\frac{2\pi \delta}{\lambda}=\frac{2}{\pi}."